# Viability of $A_{4}$, $S_{4}$ and $A_{5}$ Flavour Symmetries in Light of the First JUNO Result

### v1

@@ Introduction.
After only two months of data-taking, the JUNO experiment released its first physics result, determining the solar neutrino mixing angle, $\theta_{12}$, or more precisely $\sin^{2}\theta_{12}$, and the solar mass-squared difference, $\Delta m^{2}_{21}\equiv m_{2}^{2}-m_{1}^{2}$, to unprecedented precision [1]:
@@
In this work, we first update our previous analysis [10] employing the results of the latest global fit to neutrino oscillation data performed by the NuFIT collaboration in September 2024 [11, 12]. [^1] See [13] for an independent global analysis of neutrino oscillation data that yields similar results. We find that five (four) cases leading to sharp predictions for $\sin^{2}\theta_{12}$ are compatible with the global data at the $3\sigma$ confidence level, assuming normal (inverted) neutrino mass ordering, denoted further as NO (IO). Next, we perform a statistical analysis including the recent JUNO result, eq. (1). We find that only three (two) mixing patterns survive at $3\sigma$ after the JUNO measurement. We also discuss the predictions of these mixing patterns for the Dirac CP-violating (CPV) phase, $\delta$, present in the PMNS lepton mixing matrix.
@@
This letter is organised as follows. First, we briefly summarise the symmetry-based mixing patterns of interest. Then, after describing the procedure used in our statistical analysis, we present our results in terms of likelihood plots for $\sin^{2}\theta_{12}$ and $\cos\delta$. Finally, we summarise our findings and draw conclusions.
@@
Residual symmetries $G_{e}$ and $G_{\nu}$ of the charged lepton and neutrino mass matrices, respectively, lead to relations between (i) (some of) the mixing angles and (ii) (some of) the mixing angles and the Dirac CPV phase, $\delta$, in the PMNS matrix. As shown in [9] (see also [10]), if $G_{e}=Z_{k}$, $k>2$ or $Z_{m}\times Z_{n}$, $m,n\geq 2$ and $G_{\nu}=Z_{2}$, the solar neutrino mixing parameter $\sin^{2}\theta_{12}$ (cosine of the CPV phase, $\cos\delta$) is expressed in terms of $\theta_{13}$ ($\theta_{13}$ and $\theta_{23}$) and the parameters $\theta^{\circ}_{ij}$ fixed by the underlying symmetries. This pattern of residual symmetries, denoted as pattern B, leads to two cases, B1 and B2, depending on the plain, not fixed by $G_{\nu}=Z_{2}$, in which a $U(2)$ transformation acts. In case B1, we have
@@
$\cos\delta=-\frac{\cos^{2}\theta_{13}(\cos^{2}\theta^{\circ}_{12}\cos^{2}\theta^{\circ}_{23}-\cos^{2}\theta_{23})+\sin^{2}\theta^{\circ}_{12}(\cos^{2}\theta_{23}-\sin^{2}\theta_{13}\sin^{2}\theta_{23})}{\sin 2\theta_{23}\sin\theta_{13}|\sin\theta^{\circ}_{12}|(\cos^{2}\theta_{13}-\sin^{2}\theta^{\circ}_{12})^{\frac{1}{2}}}\,.$ (3)
@@
$\cos\delta=\frac{\cos^{2}\theta_{13}(\sin^{2}\theta^{\circ}_{12}-\cos^{2}\theta_{23})+\cos^{2}\theta^{\circ}_{12}\cos^{2}\theta^{\circ}_{13}(\cos^{2}\theta_{23}-\sin^{2}\theta_{13}\sin^{2}\theta_{23})}{\sin 2\theta_{23}\sin\theta_{13}|\cos\theta^{\circ}_{12}\cos\theta^{\circ}_{13}|(\cos^{2}\theta_{13}-\cos^{2}\theta^{\circ}_{12}\cos^{2}\theta^{\circ}_{13})^{\frac{1}{2}}}\,.$ (5)
@@
Table 1: Cases derived from $A_{4}$, $S_{4}$ and $A_{5}$ and predicting $\sin^{2}\theta_{12}$ within its $3\sigma$ range obtained by NuFIT. Case B1 is common to all the three groups. The values of $\sin^{2}\theta_{12}$ ($\cos\delta$) are obtained using the best-fit values of $\sin^{2}\theta_{13}$ ($\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$). An asterisk means that physical values of $\cos\delta$ cannot be obtained employing the best-fit values of the mixing angles.
The irrational values of $\sin^{2}\theta^{\circ}_{ij}$ appearing for the cases originating from $A_{5}$ symmetry are related to the golden ratio $r=(1+\sqrt{5})/2$ as follows: $1/(2+r)\approx 0.276$, $1/(4r^{2})\approx 0.095$, and $(3-r)/4\approx 0.345$ [9]. The quoted predictions for $\sin^{2}\theta_{12}$ and $\cos\delta$ are obtained using the best-fit values of $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ (with Super-Kamiokande atmospheric data) from NuFIT 6.0 [11, 12]. Case C5 characterised by $\sin^{2}\theta^{\circ}_{12}=1/4$ and common to $S_{4}$ and $A_{5}$ leads to $\sin^{2}\theta_{12}\approx 0.256$, a value excluded at $4.9\sigma$ by NuFIT and at $6.1\sigma$ by the first JUNO measurement. This is why this case does not appear in the Table. The mixing patterns in Table 1 have been studied in the context of current and future neutrino oscillation experiments in [10, 24, 25].
@@
We would like to mention that there exist other phenomenological sum rules for $\cos\delta$. The PMNS matrix $U_{\mathrm{PMNS}}=U_{e}^{\dagger}U_{\nu}$, with the unitary matrix $U_{\nu}$ ($U_{e}$) originating from the diagonalisation of the neutrino (charged lepton) mass term. Assuming that (i) $U_{\nu}$ has one of the five symmetry forms, namely, tri-bimaximal [28, 29, 30, 31], bimaximal [32, 33, 34], golden ratio type A [35, 36], golden ratio type B [37, 38], or hexagonal [39, 40] mixing, and (ii) $U_{e}$ is given by either a unitary rotation in one plane or a product of two unitary rotations in any two different planes, sum rules for $\cos\delta$ were derived in [41, 42]. The test of these phenomenologically derived sum rules might provide independent evidence, if any, for a non-Abelian discrete symmetry at the origin of the experimentally observed peculiar lepton mixing pattern of two large and one small mixing angles.
@@
To assess the viability of the mixing patterns in Table 1, we construct an approximate likelihood function as discussed in [10]. Namely, for the observable $\alpha$ of interest ($\sin^{2}\theta_{12}$ or $\cos\delta$ in our case), it is given by
@@
$\chi^{2}(\alpha)=\min\Bigg[\chi^{2}(\vec{x})\bigg|_{\hbox{\set@color\hskip 19.33424pt\hskip-18.22884pt\hbox{\set@color\small$\alpha=\mathrm{const}$}\hskip-18.22884pt\hskip-19.33424pt\raisebox{7.5357pt}{\hbox{\set@color\small sum rules}}\hskip-19.33424pt\hskip 19.33424pt}}\Bigg]\qquad\text{with}\qquad\vec{x}=(\sin^{2}\theta_{12},\sin^{2}\theta_{13},\sin^{2}\theta_{23},\delta)\,$, (11)
and the total $\chi^{2}(\vec{x})$ is given by the sum of the one-dimensional projections $\chi^{2}_{i}(x_{i})$:
$\chi^{2}(\vec{x})=\sum_{i=1}^{4}\chi^{2}_{i}(x_{i})\,.$ (12)
In our analysis, we first use the one-dimensional projections provided by the NuFIT collaboration on their website [12]. Next, to incorporate the JUNO result, we replace $\chi^{2}_{1}(\sin^{2}\theta_{12})$ with
@@ Results.
The top-left (top-right) panel uses as an input the NuFIT one-dimensional projections $\chi_{i}^{2}(x_{i})$ for NO (IO). The global analysis [11, 12] (with Super-Kamiokande atmospheric data) finds an overall preference for NO at the level of $2.5\sigma$ ($\Delta\chi^{2}=6.1$). However, we take a conservative approach and treat both NO and IO on equal grounds. [^2] This is the reason why the black dashed line reaches $\mathcal{L}(\sin^{2}\theta_{12})=1$ in the maximum (corresponding to $\chi^{2}_{1}(\sin^{2}\theta_{12})=0$) for both neutrino mass orderings. We note that the main differences between the NuFIT 6.0 results and those of NuFIT 3.2 used in [10] are the best-fit values of and the one-dimensional $\chi^{2}$-projections for $\sin^{2}\theta_{23}$ and $\delta$. As a consequence, case B2A5 that was found to be compatible with the global data at $3\sigma$ in [10] is now disfavoured at $3.2\sigma$ for IO, and thus, does not appear in the right panel. Further, cases B1 and B1A5 barely surpass the $3\sigma$ line. Case B2A5II is disfavoured at $3\sigma$ ($4\sigma$) for NO (IO), and thus, does not appear in the plots.
In the lower panels of Fig. 1, we implement the JUNO result in the form of eq. (13). The high-precision measurement of $\sin^{2}\theta_{12}$ results in disfavouring cases B1 (corresponding to TM2) and B1A5 at $3.8\sigma$ and $3.3\sigma$, respectively, for both NO and IO. In addition, case B2A5, while being compatible with the JUNO result at $2.6\sigma$ for NO, is disfavoured at $3.5\sigma$ for IO. Overall, out of the five (four) cases allowed by the global data at $3\sigma$ for NO (IO), only three (two) cases survive (at the same confidence level) after the JUNO measurement. The most favoured pattern, compatible with the data at $1.8\sigma$ ($1.1\sigma$) for NO (IO) is case B2S4 corresponding to TM1.
@@
For these patterns, the predicted value of $\cos\delta$ is very sensitive to the value of $\theta_{23}$, as can be understood from eqs. (3) and (5). The two maxima in $\mathcal{L}(\cos\delta)$, especially pronounced in case B2S4, are a consequence of the fact that the one-dimensional projection $\chi^{2}_{3}(\sin^{2}\theta_{23})$ has two local minima, unlike in the NuFIT 3.2 results used to obtain Fig. 3 in [10]. As we can see from the lower panels, the JUNO measurement almost does not affect the width of the likelihood profiles, that is driven by the uncertainty in the determination of the atmospheric mixing angle, $\theta_{23}$. Instead, it has an impact on the overall compatibility of the mixing patterns with the data, with only cases B2S4 and B2A5 (B2S4) surviving at the $3\sigma$ confidence level. We recall that case C9A5 does not yield a sum rule relating $\cos\delta$ to the mixing angles, and hence, it is not present in Fig. 2.
@@
Since any of the non-Abelian symmetries discussed by us are assumed to be realised at some high scale, the question arises how our results might be modified by renormalisation group (RG) running effects. The evaluation of the RG effects requires additional specific assumptions about the scale at which the symmetry is realised, the underlying model (Standard Model (SM), minimal supersymmetric SM, etc.), the nature — Dirac or Majorana — of massive neutrinos and the mechanism of neutrino mass generation (SM minimally extended with right-handed singlet neutrinos and conserved total lepton charge for Dirac neutrinos or seesaw mechanism/Weinberg dimension-5 operator for Majorana neutrinos, etc.). They further depend on the type of spectrum of neutrino masses (NO or IO), on the values of the lightest neutrino mass and, in the case of massive Majorana neutrinos, on the Majorana phases. However, in all these cases the RG effects are known to be negligible if the lightest neutrino mass is smaller than approximately 0.01 eV, see, e.g., [43, 44]. Thus, we can conclude conservatively that our results are valid at least in this case. For detailed studies of the RG effects on sum rules arising for various lepton mixing patterns, we refer the reader to [45, 46, 14].
@@
In this letter, we studied the implications of the first JUNO measurement of $\sin^{2}\theta_{12}$ [1] for lepton mixing patterns shaped by residual symmetries of the charged lepton and neutrino mass matrices, that originate from $A_{4}$, $S_{4}$ and $A_{5}$ flavour symmetries. First, we updated our previous analysis [10] of these mixing patterns using the results of the latest global fit to neutrino oscillation data performed by NuFIT [11, 12]. We found that among the mixing patterns yielding sharp predictions for $\sin^{2}\theta_{12}$, there are five (four) cases compatible with the global data at $3\sigma$ for NO (IO) (upper panels of Fig. 1). Next, we performed a statistical analysis incorporating the recent JUNO result in eq. (1) and found that the number of cases viable at $3\sigma$ is reduced to only three (two) for NO (IO) (lower panels of Fig. 1). We further derived the likelihood functions for $\cos\delta$ before and after the JUNO measurement, see Fig. 2. Our results show that the first JUNO data has strong implications for the lepton mixing patterns arising from $A_{4}$, $S_{4}$ and $A_{5}$ flavour symmetries. It is very probable that all of the remaining viable cases will be critically tested by future JUNO measurements of $\sin^{2}\theta_{12}$. However, if any of them survives the future tests, plausibly this could be case B2S4 corresponding to TM1 mixing, a sufficiently high-precision measurement of $\delta$ would provide a decisive answer on whether it is realised in nature.

### v2

@@ Introduction.
After only 59.1 days of data-taking, the JUNO experiment  [1, 2, 3, 4], conceived in [5, 6, 7, 8], released its first physics result, determining the solar neutrino mixing angle, $\theta_{12}$, or more precisely $\sin^{2}\theta_{12}$, and the solar neutrino mass-squared difference, $\Delta m^{2}_{21}\equiv m_{2}^{2}-m_{1}^{2}$, to unprecedented precision [9]:
@@
In this work, we first update our previous analysis [18] employing the results of the latest global fit to neutrino oscillation data performed by the NuFIT collaboration in September 2024 [19, 20], see Table 1.111See [21] for an independent global analysis of neutrino oscillation data that yields similar results.
Normal ordering (NO) Inverted ordering (IO)
Parameter Best-fit value $\pm 1\sigma$ $3\sigma$ range Best-fit value $\pm 1\sigma$ $3\sigma$ range
$\sin^{2}\theta_{12}$ $0.308^{+0.012}_{-0.011}$ $0.275\to 0.345$ $0.308^{+0.012}_{-0.011}$ $0.275\to 0.345$
$\sin^{2}\theta_{23}$ $0.470^{+0.017}_{-0.013}$ $0.435\to 0.585$ $0.550^{+0.012}_{-0.015}$ $0.440\to 0.584$
$\sin^{2}\theta_{13}$ $02215^{+0.00056}_{-0.00058} 0.0$$030\to 0.02388 0.022$$231^{+0.00056}_{-0.00056} 0.0$$02060\to 0.02409$
$\delta~[^{\circ}]$ $212^{+26}_{-41}$ $124\to 364$ $274^{+22}_{-25}$ $201\to 335$
Table 1: Best-fit values along with $1\sigma$ uncertainties and $3\sigma$ ranges of the lepton mixing parameters obtained in the latest global analysis of neutrino oscillation data (including Super-Kamiokande) performed by the NuFIT collaboration in September 2024 [19, 20].
We find that five (four) cases leading to sharp predictions for $\sin^{2}\theta_{12}$ are compatible with the global data at the $3\sigma$ confidence level, assuming normal (inverted) neutrino mass ordering, denoted further as NO (IO). Next, we perform a statistical analysis including the recent JUNO result, eq. (1). We find that only three (two) mixing patterns survive at $3\sigma$ after the JUNO measurement. We also discuss the predictions of these mixing patterns for the atmospheric neutrino mixing angle, $\theta_{23}$, and the Dirac CP-violating (CPV) phase, $\delta$, present in the PMNS lepton mixing matrix.
@@
This letter is organised as follows. First, we briefly summarise the symmetry-based mixing patterns of interest. Then, after describing the procedure used in our statistical analysis, we present our results in terms of likelihood plots for $\sin^{2}\theta_{12}$, $\cos\delta$ and $\sin^{2}\theta_{23}$. Finally, we summarise our findings and draw conclusions.
@@
Residual symmetries $G_{e}$ and $G_{\nu}$ of the charged lepton and neutrino mass matrices, respectively, lead to relations between (i) (some of) the mixing angles and (ii) (some of) the mixing angles and the Dirac CPV phase, $\delta$, in the PMNS matrix. As shown in [17] (see also [18]), if $G_{e}=Z_{k}$, $k>2$ or $Z_{m}\times Z_{n}$, $m,n\geq 2$ and $G_{\nu}=Z_{2}$, the solar neutrino mixing parameter $\sin^{2}\theta_{12}$ (cosine of the CPV phase, $\cos\delta$) is expressed in terms of $\theta_{13}$ ($\theta_{13}$ and $\theta_{23}$) and the parameters $\theta^{\circ}_{ij}$ fixed by the underlying symmetries. This pattern of residual symmetries, denoted as pattern B, leads to two cases, B1 and B2, depending on the plane, not fixed by the $G_{\nu}=Z_{2}$ symmetry, in which a $U(2)$ transformation acts. In case B1, we have
@@
$\cos\delta=-\frac{\cos^{2}\theta_{13}(\cos^{2}\theta^{\circ}_{12}\cos^{2}\theta^{\circ}_{23}-\cos^{2}\theta_{23})+\sin^{2}\theta^{\circ}_{12}(\cos^{2}\theta_{23}-\sin^{2}\theta_{13}\sin^{2}\theta_{23})}{\sin 2\theta_{23}\sin\theta_{13}|\sin\theta^{\circ}_{12}|(\cos^{2}\theta_{13}-\sin^{2}\theta^{\circ}_{12})^{\frac{1}{2}}}\,$, (3)
where $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ are functions of two a priori free real parameters — an angle $\theta\in[0,\pi)$ and a phase $\phi\in[0,2\pi)$ — entering the $U(2)$ transformation [17] (see also [32]):
$\sin^{2}\theta_{13}=\cos^{2}\theta^{\circ}_{12}\sin^{2}\theta\,$, (4)
$\sin^{2}\theta_{23}=\frac{\cos^{2}\theta^{\circ}_{23}\sin^{2}\theta\sin^{2}\theta^{\circ}_{12}+\cos^{2}\theta\sin^{2}\theta^{\circ}_{23}-\frac{1}{2}\sin 2\theta\sin 2\theta^{\circ}_{23}\sin\theta^{\circ}_{12}\cos\phi}{1-\sin^{2}\theta_{13}}\,.$ (5)
@@
$\cos\delta=\frac{\cos^{2}\theta_{13}(\sin^{2}\theta^{\circ}_{12}-\cos^{2}\theta_{23})+\cos^{2}\theta^{\circ}_{12}\cos^{2}\theta^{\circ}_{13}(\cos^{2}\theta_{23}-\sin^{2}\theta_{13}\sin^{2}\theta_{23})}{\sin 2\theta_{23}\sin\theta_{13}|\cos\theta^{\circ}_{12}\cos\theta^{\circ}_{13}|(\cos^{2}\theta_{13}-\cos^{2}\theta^{\circ}_{12}\cos^{2}\theta^{\circ}_{13})^{\frac{1}{2}}}\,$, (7)
where the expressions for $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ take the following form [17] (see also [32]):
$\sin^{2}\theta_{13}=\cos^{2}\theta^{\circ}_{13}\sin^{2}\theta^{\circ}_{12}\sin^{2}\theta+\sin^{2}\theta^{\circ}_{13}\cos^{2}\theta+\frac{1}{2}\sin 2\theta\sin 2\theta^{\circ}_{13}\sin\theta^{\circ}_{12}\cos\phi\,$, (8)
$\sin^{2}\theta_{23}=\frac{\cos^{2}\theta^{\circ}_{12}\sin^{2}\theta}{1-\sin^{2}\theta_{13}}\,.$ (9)
@@
Table 2: Cases derived from $A_{4}$, $S_{4}$ and $A_{5}$ and predicting $\sin^{2}\theta_{12}$ within its $3\sigma$ range obtained by NuFIT. Case B1 is common to all the three groups. The values of $\sin^{2}\theta_{12}$ ($\cos\delta$) are obtained using the best-fit values of $\sin^{2}\theta_{13}$ ($\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ for NO and IO). An asterisk means that physical values of $\cos\delta$ cannot be obtained employing the best-fit values of the mixing angles.
The irrational values of $\sin^{2}\theta^{\circ}_{ij}$ appearing for the cases originating from $A_{5}$ symmetry are related to the golden ratio $r=(1+\sqrt{5})/2$ as follows: $1/(2+r)\approx 0.276$, $1/(4r^{2})\approx 0.095$, and $(3-r)/4\approx 0.345$ [17]. The quoted predictions for $\sin^{2}\theta_{12}$ and $\cos\delta$ are obtained using the best-fit values of $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ (with Super-Kamiokande atmospheric data) from NuFIT 6.0 [19, 20], see Table 1. Case C5 characterised by $\sin^{2}\theta^{\circ}_{12}=1/4$ and common to $S_{4}$ and $A_{5}$ leads to $\sin^{2}\theta_{12}\approx 0.256$, a value excluded at $4.9\sigma$ by NuFIT and at $6.1\sigma$ by the first JUNO measurement. This is why this case does not appear in Table 2. The mixing patterns in Table 2 have been studied in the context of current and future neutrino oscillation experiments in [18, 32, 33].
@@
We would like to mention that there exist other phenomenological sum rules for $\cos\delta$. The PMNS matrix $U_{\mathrm{PMNS}}=U_{e}^{\dagger}U_{\nu}$, with the unitary matrix $U_{\nu}$ ($U_{e}$) originating from the diagonalisation of the neutrino (charged lepton) mass term. Assuming that (i) $U_{\nu}$ has one of the five symmetry forms, namely, tri-bimaximal [36, 37, 38, 39], bimaximal [40, 41, 42], [^2] Bimaximal mixing can be obtained, in particular, from $U(1)$ symmetry associated with the conservation of the non-standard lepton charge $L^{\prime}=L_{e}-L_{\mu}-L_{\tau}$ [43]. golden ratio type A [44, 45], golden ratio type B [46, 47], or hexagonal [48, 49] mixing, and (ii) $U_{e}$ is given by either a unitary rotation in one plane or a product of two unitary rotations in any two different planes, sum rules for $\cos\delta$ were derived in [50], and the respective predictions for $\cos\delta$ were analysed in detail in [51]. The test of these phenomenologically derived sum rules might provide independent evidence, if any, for a non-Abelian discrete symmetry at the origin of the experimentally observed peculiar lepton mixing pattern of two large and one small mixing angles.
@@
To assess the viability of the mixing patterns in Table 2, we construct an approximate likelihood function, similarly to the procedure discussed in [18]. Namely, for the observable $\alpha$ of interest ($\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ or $\cos\delta$), it is given by
@@
$\chi^{2}(\alpha)=\min\Bigg[\chi^{2}(\vec{x})\bigg|_{\hbox{\set@color\hskip 19.33424pt\hskip-19.33424pt\hbox{\set@color\small sum rules}\hskip-19.33424pt\hskip-18.22884pt\raisebox{8.25pt}{\hbox{\set@color\small$\alpha=\mathrm{const}$}}\hskip-18.22884pt\hskip 19.33424pt}}\Bigg$]\,, (15)
with
$\vec{x}=(\sin^{2}\theta_{12},\sin^{2}\theta_{13},\sin^{2}\theta_{23},\delta)\,$, (16)
where the theoretical relations between the standard mixing parameters $\vec{x}$ and the model parameters $(\theta,\phi)$, eqs. (2)–(5) for patterns B1, eqs. (6)–(9) for patterns B2, etc., have been taken into account. The NuFIT collaboration does not provide the total $\chi^{2}(\vec{x})$ obtained in the global analysis. However, they do provide one-dimensional projections $\chi^{2}_{i}(x_{i})$ for each mixing parameter $x_{i}$, as well as two-dimensional projections $\chi^{2}_{ij}(x_{i},x_{j})$ for each pair of the mixing parameters $(x_{i},x_{j})$, $i\neq j$, on their website [20]. We approximate the total $\chi^{2}(\vec{x})$ as the following sum: [^3] We have checked that the approximate $\chi^{2}(\vec{x})$ constructed from the four one-dimensional projections as $\chi^{2}(\vec{x})=\sum_{i=1}^{4}\chi^{2}_{i}(x_{i})$ reproduces extremely well the two-dimensional allowed regions in the planes $(x_{i},x_{j})$, $i\neq j$, obtained by NuFIT [19, 20], with the only exception being the regions in the plane $(\sin^{2}\theta_{23},\delta)$ for NO, that are somewhat overestimated. For this reason, we adopt the more accurate approximation in eq. (17).
$\chi^{2}(\vec{x})=\chi^{2}_{1}(\sin^{2}\theta_{12})+\chi^{2}_{2}(\sin^{2}\theta_{13})+\chi^{2}_{34}(\sin^{2}\theta_{23},\delta)\,.$ (17)
In this way, we account with a relatively high precision for non-trivial correlations between $\sin^{2}\theta_{23}$ and $\delta$ present in the current experimental data, all other correlations between the observables of interest being negligible. [^4] This was confirmed also in the discussions with I. Esteban, M. C. Gonzalez-Garcia and A. Marrone, who performed a detailed numerical study of the correlations and provided us with the results.
First, we perform the statistical analysis using the results of the global fit to neutrino oscillation data. Next, to incorporate the JUNO result, we replace $\chi^{2}_{1}(\sin^{2}\theta_{12})$ in eq. (17) with
@@ Results.
The top-left (top-right) panel uses as an input the NuFIT one-dimensional projections $\chi^{2}_{1}(\sin^{2}\theta_{12})$, $\chi^{2}_{2}(\sin^{2}\theta_{13})$ and two-dimensional projection $\chi^{2}_{34}(\sin^{2}\theta_{23},\delta)$ for NO (IO). The global analysis [19, 20] (with Super-Kamiokande atmospheric data) finds an overall preference for NO at the level of $2.5\sigma$ ($\Delta\chi^{2}=6.1$). However, we take a conservative approach and treat both NO and IO on equal grounds. [^5] This is the reason why the black dashed line reaches $\mathcal{L}(\sin^{2}\theta_{12})=1$ in the maximum (corresponding to $\chi^{2}_{1}(\sin^{2}\theta_{12})=0$) for both neutrino mass orderings. We note that the main differences between the NuFIT 6.0 results and those of NuFIT 3.2 used in [18] are the best-fit values of and the one- and two-dimensional $\chi^{2}$-projections for $\sin^{2}\theta_{23}$ and $\delta$. As a consequence, case B2A5 that was found to be compatible with the global data at $3\sigma$ in [18] is now disfavoured at $3.2\sigma$ for IO, and thus, does not appear in the right panel. Further, cases B1 and B1A5 barely surpass the $3\sigma$ line. Case B2A5II is disfavoured at $3\sigma$ ($4\sigma$) for NO (IO), and thus, does not appear in the plots.
In the lower panels of Fig. 1, we implement the JUNO result in the form of eq. (18). The high-precision measurement of $\sin^{2}\theta_{12}$ results in disfavouring cases B1 (corresponding to TM2) and B1A5 at $3.8\sigma$ and $3.3\sigma$, respectively, for both NO and IO. In addition, case B2A5, while being compatible with the JUNO result at $2.7\sigma$ for NO, is disfavoured at $3.5\sigma$ for IO. Overall, out of the five (four) cases allowed by the global data at $3\sigma$ for NO (IO), only three (two) cases survive at the same confidence level after the JUNO measurement. The most favoured pattern, compatible with the data at $1.8\sigma$ ($1.1\sigma$) for NO (IO) is case B2S4 corresponding to TM1.
@@
For these patterns, the predicted value of $\cos\delta$ is very sensitive to the value of $\theta_{23}$, as can be understood from eqs. (3) and (7). The two maxima in $\mathcal{L}(\cos\delta)$ for IO, especially pronounced in case B2S4, are a consequence of the fact that the one-dimensional projection $\chi^{2}_{3}(\sin^{2}\theta_{23})$ has two local minima, unlike in the NuFIT 3.2 results used to obtain Fig. 3 in [18]. As we can see from the lower panels, the JUNO measurement almost does not affect the width of the likelihood profiles, that is driven by the uncertainty in the determination of the atmospheric mixing angle, $\theta_{23}$. Instead, it has an impact on the overall compatibility of the mixing patterns with the data, with only cases B2S4 and B2A5 (B2S4) surviving at the $3\sigma$ confidence level for NO (IO). We recall that case C9A5 does not yield a sum rule relating $\cos\delta$ to the mixing angles, and hence, it is not present in Fig. 2.
For completeness, in Fig. 3 we show the predictions of patterns B for $\sin^{2}\theta_{23}$ before (upper panels) and after (lower panels) the JUNO result.
Figure 3: Upper panels: predictions for $\sin^{2}\theta_{23}$ obtained using the results of the global analysis of neutrino oscillation data performed by the NuFIT collaboration in September 2024 [19, 20]. Lower panels: the same after including the first JUNO measurement from November 2025 [9]. The black dashed line corresponds to the one-dimensional projection $\chi^{2}_{3}(\sin^{2}\theta_{23})$ obtained by NuFIT.
The corresponding likelihood profiles have been obtained by solving eq. (5) (eq. (9)) in cases B1 (B2) for $\theta$ for a given $\phi\in[0,2\pi)$, or alternatively, by solving the equation for $\phi$ for a given $\theta\in[0,\pi)$. In case B2S4 favoured by the data, the shape of the likelihood for IO follows that of the likelihood derived in the global analysis (black dashed line), while its magnitude is suppressed, mainly because of the sharp prediction for $\sin^{2}\theta_{12}$. Almost all the values of $\sin^{2}\theta_{23}$ from its experimental $3\sigma$ range are allowed at $3\sigma$ for IO, while $\sin^{2}\theta_{23}\in[0.438,0.513]$ at $3\sigma$ for NO. In case B2A5, the values of $\sin^{2}\theta_{23}\in[0.519,0.568]$ are compatible with the global data at $3\sigma$ assuming NO. Taking into account the JUNO result, the allowed range is slightly reduced to $\sin^{2}\theta_{23}\in[0.523,0.559]$. As we have already seen, case B2A5 is disfavoured at more than $3\sigma$ assuming IO. The likelihood profiles in cases B1 and B1A5 are relatively broad, especially for IO. However, both cases are disfavoured at almost $3\sigma$ by the global data, and at more than $3\sigma$ after taking into account the JUNO measurement.
@@
Since any of the non-Abelian symmetries discussed by us are assumed to be realised at some high scale, the question arises how our results might be modified by renormalisation group (RG) running effects. The evaluation of the RG effects requires additional specific assumptions about the scale at which the symmetry is realised, the underlying model (Standard Model (SM), minimal supersymmetric SM, etc.), the nature — Dirac or Majorana — of massive neutrinos and the mechanism of neutrino mass generation (SM minimally extended with right-handed singlet neutrinos and conserved total lepton charge for Dirac neutrinos, or seesaw mechanism/Weinberg dimension-5 operator for Majorana neutrinos, etc.). They further depend on the type of spectrum of neutrino masses (NO or IO), on the value of the lightest neutrino mass and, in the case of massive Majorana neutrinos, on the Majorana phases. However, the RG effects are known to be negligible for the NO mass spectrum if the lightest neutrino mass is smaller than approximately 0.01 eV; for the IO spectrum and massive Majorana neutrinos, the RG effects on $\theta_{12}$ can be large, but are known to be particularly strongly suppressed for, e.g., specific values of the Majorana phases [52] (see also, e.g., [53]). Thus, we can conclude conservatively that our results are valid at least in the indicated cases. For detailed studies of the RG effects on sum rules arising for various lepton mixing patterns, we refer the reader to [54, 55, 22].
@@
In this letter, we studied the implications of the first JUNO measurement of $\sin^{2}\theta_{12}$ [9] for lepton mixing patterns shaped by residual symmetries of the charged lepton and neutrino mass matrices, that originate from $A_{4}$, $S_{4}$ and $A_{5}$ flavour symmetries. First, we updated our previous analysis [18] of these mixing patterns using the results of the latest global fit to neutrino oscillation data performed by NuFIT [19, 20]. We found that among the mixing patterns yielding sharp predictions for $\sin^{2}\theta_{12}$, there are five (four) cases compatible with the global data at $3\sigma$ for NO (IO) (upper panels of Fig. 1). Next, we performed a statistical analysis incorporating the recent JUNO result in eq. (1) and found that the number of cases viable at $3\sigma$ is reduced to only three (two) for NO (IO) (lower panels of Fig. 1). We further derived the likelihood functions for $\cos\delta$ and $\sin^{2}\theta_{23}$ before and after the JUNO measurement, see Figs. 2 and 3, respectively. Our results show that the first JUNO data has strong implications for the lepton mixing patterns arising from $A_{4}$, $S_{4}$ and $A_{5}$ flavour symmetries. It is very probable that all of the remaining viable cases will be critically tested by future JUNO measurements of $\sin^{2}\theta_{12}$. However, if any of them survives the future tests, plausibly this could be case B2S4 corresponding to TM1 mixing, a sufficiently high-precision measurement of $\delta$ would provide a decisive answer on whether it is realised in nature.